OpenStudy (anonymous):

I need help answering this question: -3(4x + 3) + 4(6x + 1) = 43

jimthompson5910 (jim_thompson5910):

What do you get when you distribute

OpenStudy (anonymous):

-3(4x + 3) + 4(6x + 1) = 43 -12x + -9 + 24x + 4 = 43 -12x + -5 + 24x = 43 12x + -5 = 43 12x = 38 x = 3?

jimthompson5910 (jim_thompson5910):

You made a mistake on the 2nd to last line, it should be 12x = 48

OpenStudy (anonymous):

oh my god I make the stupidest mistakes.... thank you!

jimthompson5910 (jim_thompson5910):

no worries, for the most part, you know exactly what you're doing...very nice work

OpenStudy (anonymous):

i notice i often miss minimal steps in equations and problems, even though i know what i'm doing i'll miss an important part and throw off my entire answer. is there a way i can work on that?

OpenStudy (anonymous):

or a minimal part, not even an important part, it throws off the entire thing though

jimthompson5910 (jim_thompson5910):

just be very careful and make sure to check your work you can do the check like so -3(4x + 3) + 4(6x + 1) = 43 -3(4(4) + 3) + 4(6(4) + 1) = 43 ... Plug in x = 4 -3(16 + 3) + 4(24 + 1) = 43 -3(19) + 4(25) = 43 -57 + 100 = 43 43 = 43 So the equation is true, which verifies the answer

jimthompson5910 (jim_thompson5910):

You can also graph to make sure things are equivalent, but a simple check should do

OpenStudy (anonymous):

alright, will work on! thank you again!

jimthompson5910 (jim_thompson5910):

you're welcome